# New Paradigm of Adversarial Training: Breaking Inherent Trade-Off between Accuracy and Robustness via Dummy Classes

### v1

@@ 3 Experiments
Following previous works in this area, the evaluation of the proposed method is conducted on CIFAR-10, CIFAR-100 (Krizhevsky & Hinton, 2009) and Tiny-ImageNet (Li et al., 2015) datasets with the ResNet-18 (He et al., 2016) and WideResNet-28-10 (Zagoruyko & Komodakis, 2016) architectures. The three most commonly used AT baselines namely PGD-AT, TRADES, MART, and a SOTA method, Consistency-AT, are adopted as experimental benchmarks. Additionally, we involve 14 related SOTAs specifically for the trade-off problem in our comparison by simply adopting their advances reported in the original papers. These AT methods are respectively detailed in Appendix A.2 and A.3. An example algorithm of PGD-AT + DUCAT can be found in Appendix B.
3.1 Experimental Setup
Training details.
[…]
Threat models.
In the main experiments, we consider untargeted white-box attacks under $\ell_{\infty}$ norm, with the maximal perturbation budget $\epsilon=8/255$ and optimization step size $\alpha=2/255$, which is the most common threat model for examining the adversarial robustness. We additionally consider other categories of threat models in Appendix C.1, including the one adopting the targeted attack. For adversary capacity, we assume full model information is available, so it is also visible to the adversary that the DNN under DUCAT defense will output logits with length $2C$. However, the runtime projection is not a part of the computational graph of the model, and the specific correspondences between benign and dummy classes are also diverse and underlying. Specifically, it is not necessary to implement them in order such that class $k+C$ is exactly the dummy class of class $k$. In that case, the adversary is not able to know the one-to-one correspondences unless inferring them by multiple querying as in black-box attacks.
@@ 3.1 Experimental Setup
Figure 5: Comparison between the proposed DUCAT and the current SOTAs in the trade-off problem on CIFAR-10 and ResNet-18. The superscripts ##\##, \$currency-dollar\\$\$, and * in the figure refer to the integration with TRADES, MART and Consistency-AT, respectively. Here we highlight the DUCAT and DUCAT as they are the ones namely achieve the best adversarial robustness and clean accuracy among our solutions under this comparison settings. The result demonstrates the significant advancement contributed by our work.
3.2 Main Results
Effectiveness in improving benchmarks.
As a plug-and-play method, DUCAT can be easily integrated into the four experimental benchmarks. In this section, we demonstrate the improvement brought by such integration to confirm the wide effectiveness of our new paradigm and method. Specifically, we report both clean accuracy on benign test samples and adversarial robustness under three adversaries, namely PGD-10, PGD-100 (Madry et al., 2018) and Auto-Attack (Croce & Hein, 2020) with default settings and random start. All the reported results are averages of three runs, with the specific performance of each run acquired on the best checkpoint achieving the highest PGD-10 accuracy. The results on ResNet-18 in Table 1 show that DUCAT significantly improves all benchmarks in both accuracy and robustness across all datasets, demonstrating its general effectiveness and confirming the success of our new paradigm in breaking the current trade-off between clean and adversarial targets. Additional results on WideResNet-28-10 are provided by Table 3 in Appendix C.
Compared with SOTAs.
To better show the advancement contributed by our work, we additionally compare with the SOTA related works that directly aim at releasing the trade-off between accuracy and robustness. Specific for CIFAR-10 and ResNet-18 without extra training data, we investigate the advances in the trade-off problem over the past five years, including (Ding et al., 2020; Wu et al., 2020; Zhang et al., 2020; 2021; Addepalli et al., 2022b; a; Dong et al., 2022; Rade & Moosavi-Dezfooli, 2022; Yang & Xu, 2022; Kanai et al., 2023; Zhu et al., 2023; Cao et al., 2024; Gowda et al., 2024; Wu et al., 2024; Zhang et al., 2024), and make a comparison with the results reported in their original papers. Considering the fairness, for works without original results under such settings, we turn to reliable external sources including the RobustBench leaderboard (Croce et al., 2021) and other published papers such as Liu et al. (2021). The comparison result is shown in Figure 5, with more details deferred to Table 2 in Appendix A.3.
Low additional cost in efficiency.
Figure 6: DUCAT does not introduce additional time cost significantly.
As DUCAT doubles the last fully connected layer, the model complexity can accordingly increase, which indeed brings an additional price in the training efficiency. However, compared with the parameters of the whole model, the new parameters introduced are expected to be minor, and as a result, would not significantly degrade the efficiency. As shown in Figure 6, the comparison in time cost before and after integrating DUCAT into the four experimental benchmarks supports this statement, among which the maximum additional time cost is still less than 20%. This means the additional implementation of DUCAT can be lightweight, which also fits our vision of a more practical AT for real-world applications.
3.3 Ablation Study
Provided the practical simplicity of the proposed DUCAT, there are mainly four hyper-parameters that impact the trade-off between accuracy and robustness for different reasons, namely the specific epoch $t$ to start DUCAT, and the $\alpha$, $\beta_{1}$, $\beta_{2}$ in Equation (7). As $\alpha$ directly adjusts the learning preference, attaching more importance to either benign or adversarial samples, which has a predictable impact (i.e., either better accuracy or robustness) and is not specific to our method (e.g., the regularization parameter $\lambda$ in TRADES is based on a similar idea), we defer its study to Appendix C.2. While in this section, on CIFAR-10 and ResNet-18, we conduct ablation studies for the start epoch $t$ (Figure 7(a)), as well as the soft-label construction weight $\beta_{1}$ (Figure 7(b)) and $\beta_{2}$ (Figure 7(c)) respectively for benign and adversarial samples, on the trade-off between clean accuracy (i.e., right y-axis and green line) and adversarial robustness (i.e., left y-axis, along with blue and orange line respectively for PGD-10 and Auto-Attack).
(a) DUCAT start epoch $t$.
(b) Benign soft label weight $\beta_{1}$.
(c) Adversarial soft label weight $\beta_{2}$.
Figure 7: Comprehensive ablation studies of the proposed DUCAT on CIFAR-10 and ResNet-18 w.r.t. three hyper-parameters, namely $t$, $\beta_{1}$ and $\beta_{2}$. The y-axises are aligned across the subfigures.
The first important phenomenon observed is that an appropriate epoch $t$ to start integrating DUCAT matters. Specifically, when starting too early (e.g., $t=100$), the overfitting to the training adversary PGD-10 can be serious, resulting in up to 20% robust generalization gap to test-time Auto-Attack. Yet, if starting too late (e.g., $t=120$), the learning can be insufficient in the first place, with both accuracy and robustness unsatisfying. This also reveals an important potential of DUCAT to serve as an adversarial fine-turning technology that can swiftly enhance already implemented robust DNN without retraining them from scratch. More specifically, provided that the existing target model is originally built through conventional AT by an appropriate epoch (e.g., epoch $100$), we can resume the training from that epoch with DUCAT integrated. Then after a few epochs (e.g., before epoch $130$), the model can achieve a better trade-off between clean accuracy and robustness beating existing SOTAs. This makes DUCAT also valuable in updating real-world robust applications.
For another, though as expected, the accuracy (or robustness) increases while robustness (or accuracy) drops as $\beta_{1}$ (or $\beta_{2}$) approaches 1, as long as within a reasonable range such that the original and dummy labels respectively serve as the primary learning target of benign and adversarial samples, the specific weight $\beta_{1}$ and $\beta_{2}$ to construct soft labels mildly impact the performance of DUCAT, demonstrating its stability. Interestingly, the optimal range of $\beta_{2}$ (e.g., [0.9, 1]) is larger in value than $\beta_{1}$ (e.g., [0.7, 0.8]), which is probably because learning adversarial samples is still harder than benign ones after being separated as different objectives, thus need to ensure more training samples.
4 Conclusion
In this work, given the inherent trade-off between clean accuracy and robustness in AT and the undesirable recent progress on this problem, we reveal that always-failed samples widely exist in conventional AT, explicitly attributing such a trade-off to the overstrict assumption of the current AT paradigm. In response, we suggest a new AT paradigm with dummy classes to relax this assumption, and accordingly propose a plug-and-play DUCAT method, breaking the trade-off and outperforming four benchmarks plus 14 SOTAs. For future works, more advanced methods under the new AT paradigm might be a promising direction, and we hope this work could attract more attention and inspire further studies on different paths from the current one to both accurate and robust DNNs.
Appendix
Appendix A Details of Adversaries and AT Methods Involved
@@
In this section, we supplement the introduction of the adversarial attack technologies serving as adversaries in this work, including PGD and Auto-Attack. Also, we introduce more details about the AT methods involved, including four benchmarks namely PGD-AT (Madry et al., 2018), TRADES (Zhang et al., 2019), MART (Wang et al., 2020b) and Consistency-AT (Tack et al., 2022), as well as 14 SOTAs on the trade-off problem, respectively AWP (Wu et al., 2020), FAT (Zhang et al., 2020), MMA (Ding et al., 2020), GAIRAT (Zhang et al., 2021), ACAT (Addepalli et al., 2022b), DAAT (Yang & Xu, 2022), DAJAT (Addepalli et al., 2022b), HAT (Rade & Moosavi-Dezfooli, 2022), OAAT (Addepalli et al., 2022a), TE (Dong et al., 2022), RiFT (Zhu et al., 2023), SOVR (Kanai et al., 2023), ADR (Wu et al., 2024), CURE (Gowda et al., 2024), PART (Zhang et al., 2024), VFD (Cao et al., 2024).
A.1 Adversarial Robustness and Attack Technologies
@@
Adversarial robustness refers to the performance of DNN classifiers under malicious perturbations by any possible adversaries, which is commonly measured with the test accuracy under adversarial attacks (Bai et al., 2021). The concept of adversarial attack is proposed by Biggio et al. (2013) and Szegedy et al. (2014), followed by a basic write-box method, fast gradient sign method (FGSM) (Goodfellow et al., 2015), that generates the perturbation based on the gradient of the loss function w.r.t. the input sample. The projected gradient descent attack (PGD) (Madry et al., 2018) can be viewed as a variant of FGSM, which first produces the perturbation by iteratively running FGSM, and then projects it back into the $\epsilon$-ball of the input sample. Nowadays, as Auto-Attack (Croce & Hein, 2020) forms a parameter-free and user-independent ensemble of attacks, which can identify frequent pitfalls in AT practice including over-adjustment of hyper-parameters and gradient obfuscation or masking, it has been widely recognized as one of the most reliable adversaries to exam robustness.
@@ A.2 Experimental AT Benchmarks
TRADES (Zhang et al., 2019).   As another typical AT method, TRADES proposed to decompose the robust error into natural error and boundary error, thus directly balancing the trade-off between natural accuracy and robustness. Specifically, boundary error occurs when the specific data point is sufficiently close to the decision boundary that can easily cross it under slight perturbation, which is believed as one reason for the existence of adversarial samples (Bai et al., 2021; Wang et al., 2024).
@@
Consistency-AT (Tack et al., 2022).   It is found that consistency regularization forces the model to give the same output distribution when the input or weights are slightly perturbed, which fits the goal of AT when the perturbation is generated adversarially (Zhang et al., 2022). Consistency-AT proposed a new target that the predictive distributions after attacking from two different augmentations of the same instance should be similar to each other. The underlying principle is that adversarial robustness essentially refers to model stability around naturally occurring inputs, learning to satisfy such a constraint should not inherently require labels (Carmon et al., 2019), which also relaxes the current assumption from a different perspective.
@@ A.3 SOTA AT Methods on Trade-Off between Accuracy and Robustness
The trade-off between clean accuracy and adversarial robustness has been explored from various perspectives under the current AT paradigm. Below we list 14 representative SOTAs on this problem in the past five years, which are adopted to demonstrate the advancement achieved by our new AT paradigm and DUCAT method in Section 3.2. In this section, we show the sources and detailed records for these SOTAs in Table 2, which are the corresponding raw data we rely on to draw the aforementioned Figure 5. Besides, we also briefly introduce the principle for each of these SOTAs, except for FAT (Zhang et al., 2020), HAT (Rade & Moosavi-Dezfooli, 2022) and SOVR (Kanai et al., 2023) that have been introduced in Section 1.
Table 2: Corresponding to Figure 5 in Section 3.2, this table illustrates the raw data of the 14 SOTAs on the trade-off problem, demonstrating the advancement of our work.
Year Method Clean Auto-Attack Record Source
AWP (Wu et al., 2020) 85.17 47.00 Liu et al. (2021)
2020 FAT (Zhang et al., 2020) 87.97 43.90 Liu et al. (2021)
MMA (Ding et al., 2020) 85.50 37.20 original paper
2021 GAIRAT (Zhang et al., 2021) 83.22 33.35 Liu et al. (2021)
ACAT (Addepalli et al., 2022b) 82.41 49.80 original paper
DAAT (Yang & Xu, 2022) 88.31 44.32 original paper
2022 DAJAT (Addepalli et al., 2022b) 85.71 52.48 RobustBench
HAT (Rade & Moosavi-Dezfooli, 2022) 84.90 49.08 original paper
OAAT (Addepalli et al., 2022a) 80.24 51.06 RobustBench
TE (Dong et al., 2022) 83.66 49.40 original paper
2023 RiFT (Zhu et al., 2023) 83.44 53.65 original paper
SOVR (Kanai et al., 2023) 81.90 49.40 original paper
ADR (Wu et al., 2024) 82.41 50.38 original paper
2024 CURE (Gowda et al., 2024) 86.76 49.69 original paper
PART (Zhang et al., 2024) 83.77 41.41 original paper
VFD (Cao et al., 2024) 88.30 46.60 original paper
Ours (PGD-AT + DUCAT) 88.81 58.61 -
Ours (Consistency-AT + DUCAT) 89.51 57.18 -
@@
AWP (Wu et al., 2020) proposed a double perturbation mechanism that can flatten the loss landscape by weight perturbation to improve robust generalization. MMA (Ding et al., 2020) proposed to use adaptive $\epsilon$ for adversarial perturbations to directly estimate and maximize the margin between data and the decision boundary. GAIRAT (Zhang et al., 2021) proposed that a natural data point closer to (or farther from) the class boundary is less (or more) robust, and the corresponding adversarial data point should be assigned with larger (or smaller) weight. DAJAT (Addepalli et al., 2022b) aims to handle the conflicting goals of enhancing the diversity of the training dataset and training with data that is close to the test distribution by using a combination of simple and complex augmentations with separate batch normalization layers. ACAT (Addepalli et al., 2022b) is a two-step variant of DAJAT to improve computational efficiency. OAAT (Addepalli et al., 2022a) aligns the predictions of the model with that of an Oracle during AT to achieve robustness within larger bounds.
@@
Enlightening, TE (Dong et al., 2022) proposed that one-hot labels can be noisy for them because they naturally lie close to the decision boundary, which makes it essentially difficult to assign high-confident one-hot labels for all perturbed samples within the $\epsilon$-ball of them (Stutz et al., 2020; Cheng et al., 2022). So the model may try to memorize these hard samples during AT, resulting in robust overfitting. This may also be another reason why AT leads to more complicated decision boundary, and explains why robust overfitting is harder to alleviate than the clean one. Based on a similar idea, DAAT (Yang & Xu, 2022) adaptively adjusts the perturbation ball to a proper size for each of the natural examples with the help of a naturally trained calibration network.
Table 3: The additional results of integrating the proposed DUCAT to four AT benchmarks on CIFAR-10, CIFAR-100 and Tiny-ImageNet with WideResNet-28-10 under $\ell_{\infty}$ adversaries, corresponding to Table 1 in Section 3.2. All results are acquired from three runs.
Dataset Method Original +DUCAT (Ours)
   Clean PGD-10 PGD-100 Auto-Attack    Clean PGD-10 PGD-100 Auto-Attack
PGD-AT 87.49 55.81 54.37 50.96 90.02 64.53 62.76 58.77
CIFAR-10 TRADES 85.35 57.15 56.17 51.88 83.62 57.77 57.00 53.78
MART 82.78 58.47 57.47 50.89 84.64 58.26 57.07 54.64
Consistency-AT 86.90 55.71 54.41 50.83 89.86 67.89 64.17 59.44
PGD-AT 59.95 32.18 31.21 27.27 70.02 34.28 30.73 27.95
CIFAR-100 TRADES 59.51 32.51 32.34 27.71 59.70 33.77 33.38 28.55
MART 56.84 34.12 33.70 27.97 60.97 36.80 35.98 32.40
Consistency-AT 60.84 32.48 31.64 27.74 72.97 35.90 32.24 28.31
PGD-AT 47.79 23.97 23.59 20.00 62.58 25.86 23.68 18.65
Tiny-ImageNet TRADES 51.14 24.84 24.58 20.02 51.91 25.77 25.39 20.22
MART 45.57 26.21 25.92 21.07 50.71 28.48 28.64 26.15
Consistency-AT 50.12 25.05 24.42 20.64 62.77 25.62 25.01 22.09
Figure 8: Corresponding to Figure 2 in Section 2.1, under the PGD-10 adversary, similar failure patterns can be observed in the confusion matrixes respectively with the protection of the four experimental AT methods.
@@
More recently, RiFT (Zhu et al., 2023) introduces module robust criticality, a measure that evaluates the significance of a given module to model robustness under worst-case weight perturbations, to exploit the redundant capacity for robustness by fine-tuning the adversarially trained model on its non-robust-critical module. ADR (Wu et al., 2024) generates soft labels as a better guidance mechanism that accurately reflects the distribution shift under attack during AT. CURE (Gowda et al., 2024) finds that selectively updating specific layers while preserving others can substantially enhance the network’s learning capacity, and accordingly leverages a gradient prominence criterion to perform selective conservation, updating, and revision of weights. PART (Zhang et al., 2024) partially reduces $\epsilon$ for less influential pixels, guiding the model to focus more on key regions that affect its outputs. VFD (Cao et al., 2024) conducts knowledge distillation from a pre-trained model optimized towards high accuracy to guide the AT model towards generating high-quality and well-separable features by constraining the obtained features of natural and adversarial examples.
Algorithm 1 DUmmy Classes-based Adversarial Training (DUCAT)

### v2

@@ 3 Experiments
Following previous AT works, the evaluation of the proposed DUCAT is conducted on CIFAR-10, CIFAR-100 (Krizhevsky and Hinton, 2009) and Tiny-ImageNet (Li et al., 2015) datasets with ResNet-18 (He et al., 2016) and WideResNet-28-10 (Zagoruyko and Komodakis, 2016) architectures. The three most commonly used AT baselines, namely PGD-AT, TRADES, MART, and an advanced method, Cons-AT, are adopted as experimental benchmarks. Additionally, we involve 18 related SOTAs specifically for the trade-off problem in our comparison by simply adopting their advances reported in the original papers. These AT methods are respectively detailed in Appendix A.2 and A.3.
There are more details supplemented in our Appendix. An example algorithm for PGD-AT + DUCAT is provided in Appendix B. The training details and threat models are deferred to Appendix C.1 and  C.2. Beyond the main experiments in this section, we further evaluate the generalization to three other categories of threat models in Appendix D, respectively with different perturbation budget $\epsilon$, $\ell_{2}$ norm, and targeted attacks. Also, two types of secondary related works, respectively introducing synthetic data and inference-time defense for a better trade-off, are additionally compared in Appendix E. Then in Appendix F, we conduct comprehensive ablation studies on four hyper-parameters, $t$, $\alpha$, $\beta_{1}$, and $\beta_{2}$, followed by further discussions in Appendix G regarding the difference with DuRM (Wang et al., 2023a) and conventional soft-label strategies, as well as the outliers observed in TRADES + DUCAT.
Effectiveness in Improving Benchmarks.
As a plug-and-play method, DUCAT can be easily integrated into the four experimental benchmarks. Here we demonstrate the improvement brought by such integration to confirm the wide effectiveness of our new paradigm and method. Specifically, we report clean accuracy on benign test samples, adversarial robustness under three adversaries, namely PGD-10, PGD-100 (Madry et al., 2018) and AA (Croce and Hein, 2020b) with default settings and random start, and two measures for the accuracy-robustness trade-off, namely Mean and NRR, as detailed in Appendix C.3. All the reported results are averages of three runs, with the specific performance of each run acquired on the best checkpoint achieving the highest PGD-10 accuracy. The results on ResNet-18 in Table 1 show that DUCAT significantly improves all benchmarks in both accuracy and robustness across all datasets, demonstrating its general effectiveness and confirming the success of our new paradigm in releasing the current trade-off between clean and adversarial targets. Additional results on WideResNet-28-10 are provided by Table 3 in Appendix C, and a discussion about the relatively less improvement of TRADES + DUCAT is in Appendix G.3.
Figure 7: Comparison between the proposed DUCAT and 18 current SOTAs in the trade-off problem on CIFAR-10 and ResNet-18. The superscripts ##\##, \$currency-dollar\\$\$, and * in the figure refer to the integration with TRADES, MART and Cons-AT, respectively. The result demonstrates the significant advancement contributed by our work.
Advancement beyond 18 SOTAs.
To better demonstrate the advancement contributed by our work, we compare it with the SOTA related works directly aiming at the same accuracy-robustness trade-off problem. Specifically, for CIFAR-10 and ResNet-18 without extra training data, we investigate the advances achieved in the trade-off problem over the past five years, including Ding et al. (2020), Wu et al. (2020), Zhang et al. (2020), Zhang et al. (2021), Chen et al. (2021), Addepalli et al. (2022a), Addepalli et al. (2022b), Dong et al. (2022), Rade and Moosavi-Dezfooli (2022), Yang and Xu (2022), Kanai et al. (2023), Zhu et al. (2023), Cao et al. (2024),  Gowda et al. (2024), Wu et al. (2024), Zhang et al. (2024) and Wang et al. (2024b), and make a comparison with the results reported in their original papers. Considering the fairness, for works without original results under such settings, we turn to reliable external sources including the RobustBench leaderboard (Croce et al., 2021) and other published papers such as Liu et al. (2021). The comparison result is shown in Figure 7, where we highlight DUCAT with PGD-AT and Cons-AT, because they are the ones that respectively achieve the best adversarial robustness and clean accuracy among our solutions under the comparison settings. More details are deferred to Table 2 in Appendix A.3.
Figure 8: DUCAT only introduces slightly more time costs (i.e., 0.59 minutes per epoch, +12.9% on average across the four benchmarks).
Low Additional Cost in Efficiency.
As DUCAT doubles the last fully connected layer, the model complexity can accordingly increase, which indeed brings an additional price in the training efficiency. However, compared with the parameters of the whole model, the new parameters introduced are expected to be minor, and as a result, would not significantly degrade the efficiency. Figure 8 shows the time cost before and after integrating DUCAT into the four experimental benchmarks, where the additional time costs on RTX4090 are only 0.32, 0.32, 0.58 and 1.15 minutes per epoch, respectively. Considering the corresponding improvements in Table 1, we believe such costs are acceptable. This result confirms the additional implementation of DUCAT can be lightweight, which also fits our vision of a more practical AT for real-world applications.
4 Conclusion
In this work, we reveal that always-failed samples widely exist in conventional adversarial training (AT), based on which we explicitly attribute the inherent trade-off between clean accuracy and robustness to the overstrict assumption of the current AT paradigm. In response, we suggest a new AT paradigm with dummy classes to relax this assumption, and propose a plug-and-play DUCAT method accordingly, which mitigates the trade-off and outperforms 18 SOTAs by integrating into four benchmarks. For future works, developing more advanced methods under the new AT paradigm might be a promising direction (e.g., with synthetic data like in Appendix E.1 or additional regularization technologies such as weight averaging). We hope this work could attract more attention and inspire further studies on different paths from the current one, that achieve both accurate and robust DNNs.
Appendix
@@ 4 Conclusion
Question Index
As the appendix contains comprehensive details that may not be of interest to every reader, to make better use of it, we provide the following question index to help readers quickly locate the points regarding their main concerns.
  • May the adversary bypass dummy classes by their correspondence to benign ones? . . . C.2&D.3
  • Why AT with external/synthetic data is not involved in the main experiments? . . . . . . . . . . E.1
  • Is the design of our dummy class similar to DuRM, thus limiting our novelty? . . . . . . . . . G.1
  • Is the design of our two-hot soft label similar to existing ones, thus limiting our novelty? . . . G.2
  • Why is the improvement of integrating DUCAT to TRADES relatively less than others? . . . G.3
Appendix A Involved Attacks and AT Methods
@@
In this section, we supplement the introduction of the adversarial attack technologies serving as adversaries in this work, including PGD and AA. Also, we introduce more details about the AT methods involved, including four benchmarks namely PGD-AT (Madry et al., 2018), TRADES (Zhang et al., 2019), MART (Wang et al., 2020b) and Cons-AT (Tack et al., 2022), as well as 18 SOTAs on the trade-off problem, respectively AWP (Wu et al., 2020), FAT (Zhang et al., 2020), MMA (Ding et al., 2020), GAIRAT (Zhang et al., 2021), KD + SWA (Chen et al., 2021), ACAT (Addepalli et al., 2022a), DAAT (Yang and Xu, 2022), DAJAT (Addepalli et al., 2022a), HAT (Rade and Moosavi-Dezfooli, 2022), OAAT (Addepalli et al., 2022b), TE (Dong et al., 2022), RiFT (Zhu et al., 2023), SOVR (Kanai et al., 2023), ADR (Wu et al., 2024), CURE (Gowda et al., 2024), PART (Zhang et al., 2024), VFD (Cao et al., 2024) and ReBAT (Wang et al., 2024b).
A.1 Adversarial Robustness and Attack Methods
@@
With the remarkable development of DNNs in various real-world applications (Zhang et al., 2023; Qin et al., 2024; Zhou et al., 2025), certain security concerns have received increasing attention in recent years (Liang et al., 2024; Lin et al., 2025; Liang et al., 2025). Among them, adversarial robustness refers to the performance of DNN classifiers under malicious perturbations by any possible adversaries, which is commonly measured with the test accuracy under adversarial attacks (Bai et al., 2021). The concept of adversarial attack is proposed by Biggio et al. (2013) and Szegedy et al. (2014), followed by a basic write-box method, fast gradient sign method (FGSM) (Goodfellow et al., 2015), that generates the perturbation based on the gradient of the loss function w.r.t. the input sample. The projected gradient descent attack (PGD) (Madry et al., 2018) can be viewed as a variant of FGSM, which first produces the perturbation by iteratively running FGSM, and then projects it back into the $\epsilon$-ball of the input sample. Nowadays, as Auto-Attack (AA) (Croce and Hein, 2020b) forms a parameter-free and user-independent ensemble of attacks, which can identify frequent pitfalls in AT practice including over-adjustment of hyper-parameters and gradient obfuscation or masking, it has been widely recognized as one of the most reliable adversaries to exam robustness.
@@ A.2 Experimental AT Benchmarks
TRADES (Zhang et al., 2019). As another typical AT method, TRADES proposed to decompose the robust error into natural error and boundary error, thus directly balancing the trade-off between natural accuracy and robustness. Specifically, boundary error occurs when the specific data point is sufficiently close to the decision boundary that can easily cross it under slight perturbation, which is believed as one reason for the existence of adversarial samples (Bai et al., 2021; Wang et al., 2024a).
@@
Cons-AT (Tack et al., 2022). It is found that consistency regularization forces the model to give the same output distribution when the input or weights are slightly perturbed, which fits the goal of AT when the perturbation is generated adversarially (Zhang et al., 2022). Cons-AT proposed a new target that the predictive distributions after attacking from two different augmentations of the same instance should be similar to each other. The underlying principle is that adversarial robustness essentially refers to model stability around naturally occurring inputs, learning to satisfy such a constraint should not inherently require labels (Carmon et al., 2019), which also relaxes the current assumption from a different perspective.
@@ A.3 SOTA AT Methods on Trade-Off between Accuracy and Robustness
The trade-off between clean accuracy and adversarial robustness has been explored from various perspectives under the current AT paradigm. Below we list 18 representative SOTAs on this problem in the past five years, which are adopted to demonstrate the advancement achieved by our new AT paradigm and DUCAT method in Section 3. In this section, we show the sources and detailed records for these SOTAs in Table 2, which are the corresponding raw data we rely on to draw the aforementioned Figure 7. Besides, we also briefly introduce the principle for each of these SOTAs, except for FAT (Zhang et al., 2020), HAT (Rade and Moosavi-Dezfooli, 2022) and SOVR (Kanai et al., 2023) that have been introduced in Section 1.
Table 2: Corresponding to Figure 7 in Section 3, this table illustrates the raw data of the 18 SOTAs on the trade-off problem, demonstrating the advancement of our work. Specifically, the best results for each measure are marked in bold, while the sub-optimal results are marked with underlines. Note that ReBAT is the only one here with PreActResNet-18 due to the lack of results on ResNet-18 in its original paper.
Year Method Clean AA Mean NRR Record Source
AWP (Wu et al., 2020) 85.17 47.00 66.085 60.573 Liu et al. (2021)
2020 FAT (Zhang et al., 2020) 87.97 43.90 65.935 58.571 Liu et al. (2021)
MMA (Ding et al., 2020) 85.50 37.20 61.350 51.844 Original paper
2021 GAIRAT (Zhang et al., 2021) 83.22 33.35 58.285 47.618 Liu et al. (2021)
KD + SWA (Chen et al., 2021) 84.65 52.14 68.395 64.532 Original paper
ACAT (Addepalli et al., 2022a) 82.41 49.80 66.105 62.083 Original paper
DAAT (Yang and Xu, 2022) 88.31 44.32 66.315 59.020 Original paper
2022 DAJAT (Addepalli et al., 2022a) 85.71 52.48 69.095 65.100 RobustBench
HAT (Rade and Moosavi-Dezfooli, 2022) 84.90 49.08 66.990 62.202 Original paper
OAAT (Addepalli et al., 2022b) 80.24 51.06 65.650 62.408 RobustBench
TE (Dong et al., 2022) 83.66 49.40 66.530 62.119 Original paper
2023 RiFT (Zhu et al., 2023) 83.44 53.65 68.545 65.308 Original paper
SOVR (Kanai et al., 2023) 81.90 49.40 65.650 61.628 Original paper
ADR (Wu et al., 2024) 82.41 50.38 66.395 62.532 Original paper
CURE (Gowda et al., 2024) 86.76 49.69 68.225 63.190 Original paper
2024 PART (Zhang et al., 2024) 83.77 41.41 62.590 55.423 Original paper
VFD (Cao et al., 2024) 88.30 46.60 67.450 61.005 Original paper
ReBAT (Wang et al., 2024b) 78.71 51.49 65.100 62.255 Original paper
Ours PGD-AT + DUCAT 88.81 58.61 73.710 70.617 -
Cons-AT + DUCAT 89.51 57.18 73.345 69.782 -
@@
AWP (Wu et al., 2020) proposed a double perturbation mechanism that can flatten the loss landscape by weight perturbation to improve robust generalization. MMA (Ding et al., 2020) proposed to use adaptive $\epsilon$ for adversarial perturbations to directly estimate and maximize the margin between data and the decision boundary. GAIRAT (Zhang et al., 2021) proposed that a natural data point closer to (or farther from) the class boundary is less (or more) robust, and the corresponding adversarial data point should be assigned with larger (or smaller) weight. DAJAT (Addepalli et al., 2022a) aims to handle the conflicting goals of enhancing the diversity of the training dataset and training with data that is close to the test distribution by using a combination of simple and complex augmentations with separate batch normalization layers. ACAT (Addepalli et al., 2022a) is a two-step variant of DAJAT to improve computational efficiency. OAAT (Addepalli et al., 2022b) aligns the predictions of the model with that of an Oracle during AT to achieve robustness within larger bounds.
@@
Enlightening, TE (Dong et al., 2022) proposed that one-hot labels can be noisy for them because they naturally lie close to the decision boundary, which makes it essentially difficult to assign high-confident one-hot labels for all perturbed samples within the $\epsilon$-ball of them (Stutz et al., 2020; Cheng et al., 2022). So the model may try to memorize these hard samples during AT, resulting in robust overfitting. This may also be another reason why AT leads to more complicated decision boundary, and explains why robust overfitting is harder to alleviate than the clean one. Based on a similar idea, KD + SWA (Chen et al., 2021) investigates two empirical means to inject more learned smoothening during AT, namely leveraging knowledge distillation and self-training to smooth the logits, as well as performing stochastic weight averaging (Izmailov et al., 2018) to smooth the weights, and DAAT (Yang and Xu, 2022) adaptively adjusts the perturbation ball to a proper size for each of the natural examples with the help of a naturally trained calibration network.
@@
More recently, RiFT (Zhu et al., 2023) introduces module robust criticality, a measure that evaluates the significance of a given module to model robustness under worst-case weight perturbations, to exploit the redundant capacity for robustness by fine-tuning the adversarially trained model on its non-robust-critical module. ADR (Wu et al., 2024) generates soft labels as a better guidance mechanism that accurately reflects the distribution shift under attack during AT. CURE (Gowda et al., 2024) finds that selectively updating specific layers while preserving others can substantially enhance the network’s learning capacity, and accordingly leverages a gradient prominence criterion to perform selective conservation, updating, and revision of weights. PART (Zhang et al., 2024) partially reduces $\epsilon$ for less influential pixels, guiding the model to focus more on key regions that affect its outputs. VFD (Cao et al., 2024) conducts knowledge distillation from a pre-trained model optimized towards high accuracy to guide the AT model towards generating high-quality and well-separable features by constraining the obtained features of natural and adversarial examples. ReBAT (Wang et al., 2024b) views AT as a dynamic mini-max game between the model trainer and the attacker, and proposes to alleviate robust overfitting by rebalancing the two players by either regularizing the trainer’s capacity or improving the attack strength.
Appendix B DUCAT Algorithm
In this section, we provide the specific algorithm to facilitate the understanding of the proposed DUCAT. This algorithm demonstrates the basic format of DUCAT (i.e., integrated with PGD-AT), and focuses on its novel parts, thus leaving out some general training details such as optimizer settings,
Algorithm 1 DUmmy Classes-based AT (DUCAT)